OpenStudy (anonymous):

How do You turn 0.56 into a fraction

OpenStudy (anonymous):

14/25

OpenStudy (aum):

\[ 0.56 = \frac{56}{100} = \frac{28}{50} = \frac{14}{25} \]

OpenStudy (anonymous):

0.56=56/100 reduce 56/100 (56/100) / 2=28/50 (28/50) / 2 = 14/25 14/25 is as far as you can reduce